OpenStudy (technodynamic):

The normal boiling point of ethanol is 78 degrees celsius. At 41 degrees Celsius it has a vapor pressure of 169 torr. Calculate the molar heat of vaporation, Hvap, for ethanol C2H5OH. R=8.314J/(mol K). Use the Clausius-clapeyron equation. ln P2 = lnP1+(Hvap/R)(1/T1-1/T2) Please explain step-by-step, i have to learn this for a test. In return, I will reward you with a trophy.

OpenStudy (abb0t):

You're basically rearranging the clausius clapeyron equation around and solving for \(\sf \Delta H_{vap}\). This is simply a little more applied algebra, in the sense that you need to know your rules. I will start it off for you, and you can finish it yourself: \(\sf ln(P_2)-ln(P_1)=\frac{H_{vap}}{R}[\frac{1}{T}-\frac{1}{T'}] \Rightarrow \color{red}{R~\large\frac{ln(\frac{P_2}{P_1})}{}}... = \Delta H_{vap}[...]\)

OpenStudy (technodynamic):

I came up with 37.3 kJ/mol. Did I solve correctly?

OpenStudy (abb0t):

What did you use for you're second pressure?

OpenStudy (technodynamic):

760 torr

OpenStudy (abb0t):

Yes. That's correct.